# The Wavefront Set: Bounds for the Langlands parameter

### v2

@@ 1.5. Conjectures
Let ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ denote the geometric wavefront set of an admissible representation of ${\mathbf{G}}^{\omega}({\mathsf{k}})$. This is a collection of adjoint nilpotent orbits in $\mathfrak{g}(\bar{\mathsf{k}})$. The nilpotent orbits in $\mathfrak{g}(\bar{\mathsf{k}})$ are ordered with respect to the partial order given by the Zariski closure. The nilpotent orbits in ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ are by definition non-comparable in the closure order. In almost all known examples, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ consists of a single nilpotent orbit if $\pi$ is irreducible, but not always as recent examples of Tsai [Ts] have shown.
@@
While this paper was being finalized, we learned that Conjecture 1.1, as well as the connection with Conjecture 1.9 and Jiang’s conjecture, have also been formulated independently and studied for classical groups by Hazeltine, Liu, Lo, and Shahidi [HLLS].
@@
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}(\mathbb{O}^{\vee}_{\mathsf{FT}}),$ (1.3)
whenever we know (or expect) that the action of ${\mathsf{AZ}}$ on the space of geometric parameter is given by the geometric Fourier transform $\operatorname{FT}$. This is the case for unipotent representations.
@@
The conjecture aims to generalize the recent results for Lusztig unipotent representations (trivial $\varphi(I_{\mathsf{k}})$) with real infinitesimal character, see [CMBO1], and for unipotent supercuspidal representations [CMBO2]. In these cases, the upper bound is an equality.
But even in the case of unipotent representations there is an example for the split group $E_{7}$ [CMBO1, §1] that shows that for unipotent representations with nonreal infinitesimal character, the upper bound can be strict.
At the end of section 5, we will comment on other possible versions of the conjecture involving the duality $d_{S}$ and ${}^{K}\!\underline{\mathsf{WF}}(\pi)$.
@@
For every irreducible ${\mathsf{GL}}(n,{\mathsf{k}})$-representation $\pi$,
${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=d^{\vee}(\mathbb{O}_{\pi}^{\vee}),$ (1.4)
@@
To exemplify Conjecture 1.1, we will verify it for the depth-zero supercuspidal representations of classical groups, using the Langlands parametrization of Lust-Stevens [LuSt], and of $G_{2}$, using Aubert-Xu [AX].
@@
If the representation $\pi(s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ is generic, then its geometric wavefront set is the principal orbit. For the conjecture to hold, we would need $\mathbb{O}^{\vee}_{\mathsf{FT}}=0$, so ${{\mathcal{O}}}^{\vee}$ would need to be the open dense orbit in ${\mathfrak{h}}^{\vee}_{q}$. Moreover, when $H^{\vee}(s)$ is connected, ${{\mathcal{L}}}^{\vee}$ would have to be the trivial local system.
In the case of supercuspidal representations, Conjecture 1.1 takes a simpler, more concrete form. If $\pi$ is supercuspidal, the geometric Langlands parameter is expected to satisfy
@@
    ${{\mathcal{O}}}^{\vee}={{\mathcal{O}}}_{s}^{\vee}$, the unique open dense $H^{\vee}(s)$-orbit in ${\mathfrak{h}}^{\vee}_{q}$ (this is the case whenever the representation $\pi$ is tempered).
@@
    In addition, it is immediate from the definition that
    ${\mathsf{AZ}}(\pi)=\pi.$ (1.5)
@@
Let $\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{s},{{\mathcal{L}}}^{\vee})$ be an irreducible supercuspidal ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation in the L-packet $\Pi(\varphi)=\Pi(s,{{\mathcal{O}}}^{\vee}_{s})$. Set ${\mathbb{O}}^{\vee}_{s}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{s}$. Then
@@
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}({\mathbb{O}}^{\vee}_{s}).$
    Equivalently, ${\mathbb{O}}^{\vee}_{s}\leq d({}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi))$.
@@
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=d^{\vee}({\mathbb{O}}^{\vee}_{s}).$
The first part of the conjecture is clearly just a particular case of Conjecture 1.1.
@@
Notice that Conjecture 1.6 is vacuously true for the nonsingular supercuspidal representations, in the sense of Kaletha [Kal], because ${\mathbb{O}}^{\vee}_{s}=0$ so $d^{\vee}({\mathbb{O}}^{\vee}_{s})$ is the principal geometric orbit. In this case, the upper bound is attained only for the generic constituents of the L-packets. In section 3.1, we will compute the wavefront sets of the regular depth-zero supercuspidals parameterized by Kazhdan-Varshavsky [KV] and DeBacker-Reeder [DBR].
@@
Suppose ${\mathbf{G}}({\mathsf{k}})$ is a symplectic, special orthogonal, or unitary group. Conjecture 1.6 holds for all depth-zero supercuspidal representations, using the Langlands parametrization of Lust-Stevens [LuSt].
We mention that Conjectures 1.1 and 1.6 are compatible with a conjecture about Arthur packets formulated by D. Jiang. Suppose $\varphi$ is a tempered L-parameter. Then ${\mathsf{AZ}}(\Pi(\varphi))$ is expected to be an Arthur packet. Let $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$ be one of these “anti-tempered” representations. Then $\pi^{\prime}={\mathsf{AZ}}(\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{s},{{\mathcal{L}}}^{\vee}))$ for some local system ${{\mathcal{L}}}^{\vee}$. Conjecture 1.1 then says that we should have
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi^{\prime})\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{s}).$ (1.6)
The orbit $G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{s}$ is precisely the “Arthur $SL(2)$${\mathbb{O}}^{\vee}_{A,\varphi}$ attached to the Arthur packet ${\mathsf{AZ}}(\Pi(\varphi))$. In other words, the expectation is
@@
for all $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$. This is known to hold (with equality) if $\varphi$ is a “real” unramified L-parameter, see [CMBO3]. Jiang’s conjecture [Ji, Conjecture 4.2] says that this inequality holds for all Arthur packets, when they are defined (not just the above anti-tempered packets) and moreover that there exists one constituent of the Arthur packet where the equality is achieved. One can naturally reformulate Conjecture 1.1 in this case as follows:
@@
Let $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ be an irreducible representation in the tempered L-packet $\Pi(\varphi)=\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$. Then
@@
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}^{\vee}).$
@@
In section 5, we will verify Conjecture 1.9 for $G_{2}$, using the Langlands correspondences from [AX, GS1, GS2] and the wavefront set results of [JLS, LoSa].
In section 2, we will explain that Conjecture 1.9 (via the usual desiderata of Langlands correspondence) implies Conjecture 1.1.
@@
    Firstly, let us call an irreducible representation $\pi$ anti-generic if ${\mathsf{AZ}}(\pi)$ is generic. If $\pi$ is anti-generic, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$, the regular nilpotent orbit. The only way that the upper bound conjecture could hold then is if the Langlands nilpotent parameter of $\pi$ is 00. In other words, for every anti-generic representation the image of $\mathbb{C}\subset W_{\mathsf{k}}^{\prime}$ under the Langlands parameter is trivial.
    If $\pi$ is unipotent anti-generic, then it is known that it is spherical (with respect to the maximal special parahoric) and the desired condition on the Langlands parameter holds.
    If $\pi$ is supercuspidal, since ${\mathsf{AZ}}(\pi)=\pi$, the conjecture implies that every generic supercuspidal representation has trivial nilpotent Langlands parameter.
@@
    Secondly, suppose $\mathcal{O}^{\vee}=0$, so $d^{\vee}(G^{\vee}\cdot\mathcal{O}^{\vee})=\mathcal{O}_{\textup{reg}}.$ Then Conjecture 1.9(1) is trivially satisfied, but (2) says that there should exist a representation $\pi$ in the tempered L-packet such that ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$. In other words, every non-singular tempered L-packet should contain an anti-generic representation.
Acknowledgements. This paper builds on the previous work of the first-named author with Lucas Mason-Brown and Emile Okada. It is a pleasure to thank them for many useful conversations about the wavefront set. The second-named author would like to thank George Lusztig and Cheng-Chiang Tsai for insightful discussions. We also thank Baiying Liu and Freydoon Shahidi for sharing their work with us and for an interesting correspondence. D.C. was partially supported by the EPSRC grant “New Horizons” EP/V046713/1, 1/2021-12/2023.
@@ 2. Parabolic induction
If $\pi=\operatorname{Ind}_{\mathbf{P}({\mathsf{k}})}^{{\mathbf{G}}({\mathsf{k}})}(\pi_{M})$ is a parabolically induced representation, where $\pi_{M}$ is an admissible representation of $M$, then the relation between the wavefront sets is well known, see Waldspurger [Wa]:
@@
where $\mathsf{ind}_{\mathbf{M}}^{{\mathbf{G}}}$ denotes the Lusztig-Spaltenstein induction of nilpotent orbits.
@@
Now assuming Conjecture 1.9 for tempered modules, we have ${}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\sigma))\leq d^{\vee}_{M^{\vee}}({\mathbb{O}}^{\vee}_{\sigma})$. Then Lemma 2.1(1),(3) implies
@@ 3. Depth-zero supercuspidal representations
The finite analogue of the wavefront set for $\sigma$ is the Kawanaka wavefront set $\mathsf{WF}(\sigma)$, [Kaw].
Theorem 3.1 ([CMBO2, Theorem 3.0.4]).
The canonical unramified wavefront set of $\pi(x,\sigma)$ is
${}^{K}\!\underline{\mathsf{WF}}(\pi(x,\sigma))=(\mathsf{WF}(\sigma),x).$
Lusztig gave an explicit description of $\mathsf{WF}(\mu)$, $\mu$ an irreducible representation of ${\mathbf{G}}^{\omega}_{x}({\mathbb{F}}_{q})$ in terms of his parametrization [Lu1, Theorem 4.23]. Let $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$ be the dual finite reductive group. Then $\mu$ is parametrized by a triple $(\tau,u_{0}^{\vee},y)$, where $\tau$ is a semisimple element in $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$, $u_{0}^{\vee}$ is a special unipotent element in the centralizer $C_{({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})}(\tau)$ and $y$ is an element of the unipotent family $C_{u_{0}^{\vee}}$ parametrized by $u_{0}^{\vee}$. The element $y$ will not be important for us. Let $\widetilde{C}_{u_{0}^{\vee}}=C_{\widetilde{u}_{0}^{\vee}}$ be the family obtained by applying Alvis-Curtis duality to $C_{u_{0}^{\vee}}$. Let $d_{S}^{\vee}$ be the duality map in this finite field setting, this is defined exactly as above, see [Lu1, §13.3].
@@
This means that the canonical unramified wavefront set of a supercuspidal representation $\pi(x,\sigma)$ is
${}^{K}\!\underline{\mathsf{WF}}(\pi(x,\sigma))=(d_{S}^{\vee}(\tau u_{0}^{\vee}),x).$ (3.1)
@@
$G\cdot d_{S}^{\vee}(\tau u_{0}^{\vee})\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}_{s}^{\vee}).$ (3.2)

### v3

@@ 1.5. Conjectures
Let $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ denote the geometric wavefront set of an admissible representation of ${\mathbf{G}}^{\omega}({\mathsf{k}})$. This is a collection of adjoint nilpotent orbits in $\mathfrak{g}(\bar{\mathsf{k}})$. The nilpotent orbits in $\mathfrak{g}(\bar{\mathsf{k}})$ are ordered with respect to the partial order given by the Zariski closure. The nilpotent orbits in $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ are by definition non-comparable in the closure order. In almost all known examples, $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ consists of a single nilpotent orbit if $\pi$ is irreducible, but not always, as recent examples of Tsai [Ts] have shown.
@@
While this paper was being finalized, we learned that Conjecture 1.1, as well as the connection with Conjecture 1.9 and Jiang’s conjecture, have also been formulated independently and studied for classical groups by Hazeltine, Liu, Lo, and Shahidi [HLLS]. In loc. cit., the conjectures are stated and investigated for general groups (rather than ‘inner to split’), see [HLLS, Conjecture 1.1, Remarks 1.9, 7.11].
@@
$^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}(\mathbb{O}^{\vee}_{\mathsf{FT}})=d^{\vee}(\mathbb{O}^{\vee}_{{\mathsf{AZ}}}),$ (1.4)
whenever we know (or expect) that the action of $\mathsf{AZ}$} on the space of geometric parameter is given by the geometric Fourier transform $\operatorname{FT}$.
@@
The conjecture aims to generalize the recent results for Lusztig unipotent representations (trivial $\varphi(I_{\mathsf{k}})$) with real infinitesimal character, see [CMBO1, CMBO2], and for unipotent supercuspidal representations [CMBO3]. In these cases, the upper bound is an equality.
But even in the case of unipotent representations there is an example for the split group $E_{7}$ [CMBO2, §1] that shows that for unipotent representations with nonreal infinitesimal character, the upper bound can be strict. At the end of section 5, we will comment on other possible versions of the conjecture involving the duality $d_{S}$ and $^{K}\!\underline{\mathsf{WF}}(\pi)$.
@@
For every irreducible ${\mathsf{GL}}(n,{\mathsf{k}})$-representation $\pi$, $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))$ is a singleton and
$^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\{d^{\vee}(\mathbb{O}_{\pi}^{\vee})\},$ (1.5)
@@
To exemplify Conjecture 1.1, we will verify it for the depth-zero supercuspidal representations of classical groups, using the Langlands parametrization of Lust–Stevens [LuSt], and for the irreducible representations of $G_{2}$, using Aubert–Xu [AX] and Gan–Savin [GS1, GS2].
@@
If an irreducible smooth representation $\pi$ is generic, then its geometric wavefront set is the principal orbit. For the conjecture to hold, we would need $\mathbb{O}^{\vee}_{\mathsf{AZ}}=0$, so $\mathcal{O}$$^{\vee}$ would need to be the open dense orbit in ${\mathfrak{h}}^{\vee}_{q}$. Moreover, $\mathcal{L}$$^{\vee}$ would have to be the trivial local system.
In the case of supercuspidal representations, Conjecture 1.1 takes a simpler, more concrete form. If $\pi$ is supercuspidal, the Langlands parameter is expected to satisfy
@@
    ${{\mathcal{O}}}^{\vee}={{\mathcal{O}}}_{\lambda,s}^{\vee}$, the unique open dense $H^{\vee}(s)$-orbit in ${\mathfrak{h}}^{\vee}_{q}$ (this is in fact the case whenever the representation $\pi$ is tempered). In addition,
@@
    ${\mathsf{AZ}}(\pi)=\pi$ (immediate from the definition of $\mathsf{AZ}$}).
@@
Let $\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{\lambda,s},{{\mathcal{L}}}^{\vee})$ be an irreducible supercuspidal ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation in the $L$-packet $\Pi(\varphi)=\Pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{\lambda,s})$. Set ${\mathbb{O}}^{\vee}_{\varphi}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\lambda,s}$. Then
@@
    $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}({\mathbb{O}}^{\vee}_{\varphi}).$
    Equivalently, ${\mathbb{O}}^{\vee}_{\varphi}\leq d({}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi))$.
@@
    $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\{d^{\vee}({\mathbb{O}}^{\vee}_{\varphi})\}.$
The first part of the conjecture is just a particular case of Conjecture 1.1.
@@
Notice that Conjecture 1.6 is vacuously true for the nonsingular supercuspidal representations, in the sense of Kaletha [Kal], because ${\mathbb{O}}^{\vee}_{\varphi}=0$ so $d^{\vee}({\mathbb{O}}^{\vee}_{\varphi})$ is the principal geometric orbit. In this case, the upper bound is attained only for the generic constituents of the $L$-packets. In section 3.1, we will compute the wavefront sets of the regular depth-zero supercuspidals parameterized by Kazhdan-Varshavsky [KV] and DeBacker-Reeder [DBR].
@@
Suppose ${\mathbf{G}}({\mathsf{k}})$ is a symplectic, special orthogonal, or unitary group. Conjecture 1.6 holds for all depth-zero supercuspidal representations, using the Langlands parametrization of Lust–Stevens [LuSt].
We mention that Conjectures 1.1 and 1.6 are compatible with a conjecture about Arthur packets formulated by D. Jiang. Suppose $\varphi$ is a tempered $L$-parameter. Then ${\mathsf{AZ}}(\Pi(\varphi))$ is expected to be an Arthur packet. Let $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$ be one of these “anti-tempered” representations. Then $\pi^{\prime}={\mathsf{AZ}}(\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{\lambda,s},{{\mathcal{L}}}^{\vee}))$ for some local system $\mathcal{L}$$^{\vee}$. Conjecture 1.1 then says that we should have
$^{\bar{\mathsf{k}}}\mathsf{WF}(\pi^{\prime})\leq d^{\vee}(\mathbb{O}^{\vee}_{\varphi}).$ (1.6)
[…]
@@
for all $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$. This is known to hold (with equality) if $\varphi$ is a “real” unramified $L$-parameter, see [CMBO4]. Jiang’s conjecture [Ji, Conjecture 4.2] says that this inequality holds for all Arthur packets, when they are defined (not just the above anti-tempered packets) and moreover that there exists at least one constituent of the Arthur packet where the equality is achieved. One can naturally reformulate Conjecture 1.1 in this case as follows:
@@
Let $\pi=\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{\lambda,s},{{\mathcal{L}}}^{\vee})$ be an irreducible representation in the tempered $L$-packet $\Pi(\varphi)=\Pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{\lambda,s})$. Then
@@
    $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}(\mathbb{O}^{\vee}_{\varphi}).$
@@
In section 2, we will explain that Conjecture 1.9(1) (via the usual desiderata of Langlands correspondence) implies Conjecture 1.1.
In section 5, we will verify Conjecture 1.9 for $G_{2}$, using the Langlands correspondences from [AX, GS1, GS2] and the wavefront set results of [JLS, LoSa]:
Theorem 1.10.
Conjectures 1.1 and 1.9 hold for $G_{2}$.
@@
    Firstly, let us call an irreducible representation $\pi$ anti-generic if ${\mathsf{AZ}}(\pi)$ is generic. If $\pi$ is anti-generic, then $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$, the regular nilpotent orbit. The only way that the upper bound conjecture could hold then is if the Langlands nilpotent parameter of $\pi$ is 0. In other words, for every anti-generic representation the image of $\mathbb{C}\subset W_{\mathsf{k}}^{\prime}$ under the Langlands parameter is trivial.
    If $\pi$ is unipotent anti-generic, then it is known that it is spherical (with respect to the maximal special parahoric) and the desired condition on the Langlands parameter holds. If $\pi$ is supercuspidal, since ${\mathsf{AZ}}(\pi)=\pi$, the conjecture implies that every generic supercuspidal representation has trivial nilpotent Langlands parameter.
    For classical groups, the algorithm for computing the $\mathsf{AZ}$}-dual in [AM] implies that the Langlands nilpotent parameter of an anti-generic representation is indeed trivial.
@@
    Secondly, suppose $\mathcal{O}^{\vee}=0$, so $d^{\vee}(G^{\vee}\cdot\mathcal{O}^{\vee})=\mathcal{O}_{\textup{reg}}.$ Then Conjecture 1.9(1) is trivially satisfied, but (2) says that there should exist a representation $\pi$ in the tempered $L$-packet such that $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$. In other words, every tempered $L$-packet with trivial Langlands nilpotent parameter should contain an anti-generic representation.
    This is equivalent with a conjecture of Shahidi as follows. If $\varphi\leftrightarrow(\lambda,s,{{\mathcal{O}}}^{\vee}_{\lambda,s})$ is a tempered Langlands parameter, since ${{\mathcal{O}}}^{\vee}_{\lambda,s}$ must be dense in ${\mathfrak{h}}^{\vee}_{q}$, we can only have ${{\mathcal{O}}}^{\vee}_{\lambda,s}=0$ if ${\mathfrak{h}}^{\vee}_{q}=0$. But then $\varphi$ is the only $L$-parameter for the pair $(\lambda,s)$ and since it is expected that $\mathsf{AZ}$} preserves the part $(\lambda,s)$ of the Langlands parameter, it follows that
    ${\mathsf{AZ}}(\Pi(\varphi))=\Pi(\varphi).$
@@ 2. Parabolic induction
If $\pi=i_{\mathbf{P}({\mathsf{k}})}^{{\mathbf{G}}({\mathsf{k}})}(\pi_{M})$ is a parabolically induced representation, where $\pi_{M}$ is an admissible representation of $M$, then the relation between the wavefront sets is well known, see Waldspurger [Wa]:
@@
where $\mathsf{ind}_{\mathbf{M}}^{{\mathbf{G}}}$ denotes the Lusztig–Spaltenstein induction of nilpotent orbits.
@@
Now assuming Conjecture 1.9 for tempered representations, we have $^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\sigma))\leq d^{\vee}_{M^{\vee}}({\mathbb{O}}^{\vee}_{\sigma})$. Then Lemma 2.1(1),(3) implies
@@ 3. Depth-zero supercuspidal representations
The finite analogue of the wavefront set for $\sigma$ is the Kawanaka wavefront set $\mathsf{WF}_{q}(\sigma)$, [Kaw].
Theorem 3.1 ([CMBO3, Theorem 3.0.4]).
[…]
$^{K}\!\underline{\mathsf{WF}}(\pi(x,\widetilde{\sigma}))=(\mathsf{WF}(\sigma),x).$
Lusztig gave an explicit description of $\mathsf{WF}(\mu)$, $\mu$ an irreducible representation of ${\mathcal{G}}^{\omega}_{x}$ in terms of his parametrization [Lu1, Theorem 4.23]. Let $({\mathcal{G}}^{\omega}_{x})^{\vee}$ be the dual finite reductive group. Then $\mu$ is parametrized by a triple $(\tau,u_{0}^{\vee},y)$, where $\tau$ is a semisimple element in $({\mathcal{G}}^{\omega}_{x})^{\vee}$, $u_{0}^{\vee}$ is a special unipotent element in the centralizer $C_{({\mathcal{G}}^{\omega}_{x})^{\vee}}(\tau)$ and $y$ is an element of the unipotent family $C_{u_{0}^{\vee}}$ parametrized by $u_{0}^{\vee}$. The element $y$ will not be important for us. Let $\widetilde{C}_{u_{0}^{\vee}}=C_{\widetilde{u}_{0}^{\vee}}$ be the family obtained by applying Alvis-Curtis duality to $C_{u_{0}^{\vee}}$. Let $d_{S}^{\vee}$ be the duality map in this finite field setting, this is defined exactly as above, see [Lu1, §13.3].
@@
[…]
$^{K}\!\underline{\mathsf{WF}}(\pi(x,\widetilde{\sigma}))=(d_{S}^{\vee}(\tau u_{0}^{\vee}),x).$ (3.1)
@@
$G\cdot d_{S}^{\vee}(\tau u_{0}^{\vee})\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}_{\lambda,s}^{\vee}).$ (3.2)